if 7is added to 7 and dived by 7 subtract it with 7 and multiply it by 7..............................

Grade:6

4 Answers

Arun
6 years ago
Dear Neelesh
I didn’t get your question. Please make it clear.
I will be happy to help you.
Regards
Arun (askIITians forum expert)
Chandrakanth Almale
6 years ago
7+7÷7×7-7. Following BODMAS= 7+1-7×7=7+1-49=8-49=-41 Reading the question as it is been written by studentShaumik mahapatraRajhans vidyalaya Mumbai.
Subhash Kumar Mehta
6 years ago
as by reading your que the process to solve it will be as follows >>(((7+7)/7)-7)*7>>((14/7)-7)*7 >>(2-7)*7 >>-35
Soumendu Majumdar
6 years ago
Dear Student,
According to what you have written 7 +7 = 14 divided by 7 = 2 – 7 = – 5 x 7 = – 35 
But if the question is like 
7 + 7/7 – 7 x 7 then use BODMAS rule which means division then multiplication then addition and lastly subtraction
so we get 7 + 1 – 49 = – 41
 
Hope it helps!